For what value of b is the function 𝑓(π‘₯) = { π‘₯ 2 βˆ’ 1 π‘₯ < 3 2𝑏π‘₯ π‘₯ β‰₯ 3 continuous everywhere


1
Expert's answer
2021-06-07T16:49:47-0400
"\lim\limits_{x\to3^-}f(x)=\lim\limits_{x\to3^-}(x^2-1)=(3)^2-1=8"

"\lim\limits_{x\to3^+}f(x)=\lim\limits_{x\to3^+}(2bx)=2b(3)=6b"

"\lim\limits_{x\to3^-}f(x)=8=\lim\limits_{x\to3^+}f(x)=6b"

"b=\dfrac{4}{3}"

Then


"\lim\limits_{x\to3^-}f(x)=\lim\limits_{x\to3^+}f(x)=\lim\limits_{x\to3}f(x)=8"

"f(3)=2(\dfrac{4}{3})(3)=9"

Then


"\lim\limits_{x\to3}f(x)=8=f(3)"

The function "f(x)" is continuous everywhere if "b=\dfrac{4}{3} ."